OpenStudy (anonymous):

Choose the correct graph of the given system of equations. y - 2x = -1 x + 3y = 4 graph of two lines, one with a positive slope and one with a negative slope, that intersect at the point (negative 1, 1) graph of two lines that intersect at the point (1, 1) graph of two parallel lines with positive slopes None of the above

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

it intersects at(0,2)

OpenStudy (anonymous):

it intersects at point 1,1 explanation is you get the slope intercept formula for each one then do 2x-1=-1/3x+4/3 as y=y reflexive property, then you find x is 1 then plug it back into the equation y=2x-1 and 2-1=1 so x=1, y=1.